OpenStudy (anonymous):

ali on a motorcycle with a speed of 40 km / hour.after 30 minutes followed by budi with a speed of 50 km / hour.how minutes time is needed by the budi to follow ali?

OpenStudy (anonymous):

2 hrs

OpenStudy (anonymous):

how to get it?

OpenStudy (anonymous):

take the relative velocity between two its 10km/hr since aliis followed by budi .5hours later thus by then ali must have traversed .5*40=20km.now with a relative speed of 10kmhr budi need 20hrs to catch up thank u..

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

still whats that u dont understand

OpenStudy (anonymous):

from what 10 ho km/hour and 5 hour then 20 hour

OpenStudy (anonymous):

do u understand relative velocity

OpenStudy (anonymous):

no, sorry

OpenStudy (anonymous):

ali on a motorcycle with a speed of 40 km / hour.after 30 minutes followed by budi with a speed of 50 km / hour.how minutes time is needed by the budi to follow ali? Budi: tB = (t-0.5) hour tA = t So using SA = SB; SA = SB vA x tA = vB x tB vA x t = vB x (t - 0.5) 40t = 50(t -0.5) 40t = 50t -25 50t -40t = 25 10t = 25 t = 2.5 hour or t = 2.5 x 60 minutes = 150 min do you understand @mr.tobby ???